OpenStudy (anonymous):

If f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (anonymous):

f(x) is the same as f(3) so all you do is substitute and f(3)=x+4

OpenStudy (anonymous):

is it the first one?

OpenStudy (anonymous):

It would be -1 because if x+4=3 you subtract 4 from both sides leaving you with x=-1

OpenStudy (anonymous):

f(x)=x+4 now you have to find f(x) for x=3 well here the value of x is 3 so just put x =3 f(3)=3+4=7

OpenStudy (anonymous):

its 7 or -1?

OpenStudy (anonymous):

you plug 3 in for x in your equation. 3+4=?

OpenStudy (anonymous):

Answe is second

OpenStudy (anonymous):

thnks guyz the answer was=7